OpenStudy (anonymous):

-2(3a-b)=c solve for b

OpenStudy (anonymous):

i would distribute the 2 inside the parenthese giving me: -6a+2b=c, then add 6a to both sides, which will result in : 2b=c+6a. The divivide both sides by 2, giving me answer of: b=(c+6a)/(2)